OpenStudy (anonymous):

what is 4X^3+8X^2-96X in factored form? help please!!!!

OpenStudy (anonymous):

So you have \[4x ^{3}+8x ^{2}-96x\]Hint: you can take out a 4x to make it more easier to factor.\[4x(x ^{2}+2x-24)\]

OpenStudy (anonymous):

distribute properties?

OpenStudy (anonymous):

Well, you can 'distribute' out the 4x to make it more easier to factor. How do you factor just\[x ^{2}+2x-24\] Once you found the two zeros (x) Then when finding 4x just do 4x=0

OpenStudy (anonymous):

ohh ok i get it so i factor 24

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

i got X^2+4X+6X-24?

OpenStudy (anonymous):

How did you get 4x+6x. What I did was take out a 4x from each term of the orginal problem. So \[\frac{ 4x ^{3} }{ 4x }=x ^{2} \frac{ 8x ^{2} }{ 4x }=2x \frac{ -96x }{ 4x }=-24\]So finally you would have\[4x(x ^{2}+2x-24)\]

OpenStudy (anonymous):

i found the factors (-4,6) of -24 that when u add them together (-4 + 6=+2) >> you'll end up getting +2 then i plug in the factors like this X^2-4x+6X-24

OpenStudy (anonymous):

Ok but asks you to factor so you got the factors of (-4,6) So put it into factor form of (x+a)(x+b)

OpenStudy (anonymous):

from x^2+2x-24

OpenStudy (anonymous):

is it (x-4)(x+6)?